# Factoriality of groupoid von Neumann algebras

### v2

@@ 3 Basis of a groupoid, Fourier decomposition, and the $j$ map
Since $a$ is central, $\lambda_{B}a\lambda_{B^{-1}}=\lambda_{t(B)}a$. Looking at the base coordinate $C$, this in particular means that
@@
hence we get that $a^{BCB^{-1}}(t(ghg^{-1})=a^{C}(t(h))$. Now it suffices to note that whenever an element $k\in\mathcal{G}$ lies in some bisection $E$, one has that $j(a)(k)=a^{E}\ast 1_{E}(k)=a^{E}(t(k))$. ∎
@@ 4 Factoriality of the untwisted groupoid von Neumann algebra
Let $(\mathcal{G},\mu)$ be an ergodic measured groupoid and let $A\subset\operatorname{Iso}(\mathcal{G})$ be a Borel subset with $\mu$-finite conjugacy class. Then there exist $k\in\mathbb{N}$ and finitely many mutually disjoint global bisections $V_{1},...,V_{k}$ such that $\Omega_{A}=\bigsqcup_{i}V_{i}\mod\mu_{s}$. In particular, $\mu_{s}(\Omega_{A})=k$.
@@
Let $q\colon\mathcal{R}_{\mathcal{G}}\rightarrow\mathcal{G}$ be a Borel section for the projection map $g\mapsto(t(g),s(g))$. Take an arbitrary $\varphi\in[\mathcal{R}]$ and define $W=\{q((\varphi(x),x))\mid x\in\mathcal{G}^{(0)}\}$. Then $W\in[\mathcal{G}]$ and we have that $t(g)=\varphi(s(g))$ for every $g\in W$. Observe that
@@
By ergodicty we get that for each $k$ either $\mu(E_{k})=0$ or $\mu(E_{k})=1$. Since $\mathcal{G}^{(0)}=\bigsqcup_{k}E_{k}$, we find a $k$ for which $\mu(E_{k})=1$. Then, by repeatedly applying the Lusin–Novikov theorem we find global bisections $V_{1},...,V_{k}\subset\operatorname{Iso}(\mathcal{G})$ such that $\Omega_{A}=\bigsqcup_{i}V_{i}\mod\mu_{s}$. ∎
@@
Removing the ergodicity assumption in Lemma 4.4, it is straightforward to come up with a counterexample. The transposition $(1,2)$ has conjugacy class of size (${n\choose 2}$ in the symmetric group $\mathcal{S}_{n}$. Then writing $\rho=\sum_{n=2}^{\infty}2^{-n}{n\choose 2}^{-1}<+\infty$ we get that $\mu(n)=\rho^{-1}2^{-n}{n\choose 2}^{-1}$ defines a probability measure $\mu$ on $\mathbb{N}_{\geq 2}$. Define $\mathcal{G}=\{(n,\sigma)\mid n\geq 2\text{ and }\sigma\in\mathcal{S}_{n}\}$ with unit space $(\mathbb{N}_{\geq 2},\mu)$ and source and target map given by $(n,\sigma)\mapsto n$. The conjugacy class $\Omega_{A}$ of the bisection $A=\bigcup_{n\geq 2}\{(n,(1,2))\}$ satisfies $\mu(\Omega_{A})<+\infty$ by construction, but it cannot be written as the union of finitely many bisections.
@@
Now assume condition 2 fails, so there is an element $a\in\mathcal{Z}(L(\mathcal{G}))\setminus L^{\infty}(\mathcal{G}^{(0)},\mu)$. Note that $j(a)\neq 0$ and we can assume that $j(a)|_{\mathcal{G}^{(0)}}=0$. Write
@@ 5.1 Definitions and basic properties
We say two cocycles $\omega_{1}$ and $\omega_{2}$ are cohomologous if there is a Borel map $\rho:\mathcal{G}\to\mathbb{T}$ such that $\omega_{1}(x,y)=\rho(x)\rho(y)\overline{\rho(xy)}\omega_{2}(x,y)$ for almost every $(x,y)\in\mathcal{G}^{(2)}$. Every cocycle is cohomologous to a normalized cocycle which moreover satisfies $\omega(x,x^{-1})=1$ for all $x\in\mathcal{G}$. To see this, suppose $\omega$ is a normalized cocycle and consider
@@
Now we proceed analogously as in the untwisted case by defining a ‘projective left regular representation’ in the following way: for any Borel bisection $A\subset\mathcal{G}$ we define the operator $\lambda_{A}^{\omega}\in B(L^{2}(\mathcal{G},\mu_{s})$ by
$\lambda_{A}^{\omega}1_{E}(g)=\omega(g_{1},g_{2})1_{A\cdot E}(g)\qquad\text{where }g_{1}\in A,g_{2}\in E\text{ such that }g_{1}g_{2}=g.$
The (left) twisted groupoid von Neumann algebra is given by $L_{\omega}(\mathcal{G}):=\{\lambda_{A}^{\omega}\ |\ A\subset\mathcal{G}\text{ Borel bisection}\}^{\prime\prime}$. Note that, as in the untwisted case, $L_{\omega}(\mathcal{G})$ includes a copy of $L^{\infty}(\mathcal{G}^{(0)},\mu)$ as a regular subalgebra. For any Borel bisection $A$ one has $(\lambda_{A}^{\omega})^{*}=\lambda_{A^{-1}}^{\omega}$ and one can check that $\lambda_{A}^{\omega}\lambda_{B}^{\omega}=\omega(A,B)\lambda_{A\cdot B}^{\omega}$ for any two Borel bisections $A,B\subset\mathcal{G}$.
@@
$\rho_{A}^{\omega}1_{E}(g)=\omega(g_{2},g_{1})1_{E\cdot A^{-1}}(g)\qquad\text{where }g_{1}\in A^{-1},g_{2}\in E\text{ such that }g_{2}g_{1}=g.$ (12)
One denotes the right twisted groupoid von Neumann algebra by $R_{\omega}(\mathcal{G}):=\{\rho_{A}^{\omega}\ |\ A\subset\mathcal{G}\text{ is a Borel bisection}\}^{\prime\prime}$. We again have the relations $(\rho_{A}^{\omega})^{*}=\rho_{A^{-1}}^{\omega}$ and $\rho_{A}^{\omega}\rho_{B}^{\omega}=\omega(A,B)\rho_{A\cdot B}^{\omega}$. We now show that the left $\omega$ and the right $\overline{\omega}$ representations are commutants of each other. For $f,g\in L^{2}(\mathcal{G},\mu_{s})$ we define the twisted convolution product $\ast_{\omega}$ by
@@
We say a function $f\in L^{2}(\mathcal{G})$ is a left convolver for $(\mathcal{G},\omega)$ if $f\ast_{\omega}g\in L^{2}(\mathcal{G},\mu_{s})$ for every $g\in L^{2}(\mathcal{G},\mu_{s})$. In that case if follows that $L_{f}^{\omega}(g)=f\ast_{\omega}g$ defines a bounded operator on $L^{2}(\mathcal{G},\mu_{s})$. We denote by $LC_{\omega}(\mathcal{G})$ the set consisting of all $L_{f}^{\omega}$ for left convolvers $f$. Similarly, one defines the right convolvers $RC_{\omega}(\mathcal{G})$.
@@
In particular as $f_{\mathcal{F}}\to f$ and $f_{c}F^{*}\to f^{*}$ in $L^{2}$-norm, the net $(\lambda_{f_{\mathcal{F}}}^{\omega})_{\mathcal{F}\in\mathscr{F}}$ is Cauchy. Since these elements belong in a fixed closed ball, $\lambda_{f_{\mathcal{F}}}^{\omega}$ converges in the $\sigma$-strong* operator topology to some $T\in L_{\omega}(\mathcal{G})$. I claim that $T=L_{f}^{\omega}$. For this, it suffices to show that these operators agree on bisections. For a bisection $B$, since the $\sigma$-strong* operator topology is stronger than the strong operator topology, we have the limit $T(1_{B})=\lim_{\mathcal{F}}\lambda_{f_{\mathcal{F}}}^{\omega}(1_{B})$, and
@@ 5.2 Fourier decomposition and the $j$ map again
still defines an injective linear map sending $\lambda_{A}^{\omega}$ to $1_{A}$ whenever $A\subset\mathcal{G}$ is a Borel bisection. We again consider the faithful normal state $\phi_{\mu}$ on $L_{\omega}(\mathcal{G})$ uniquely determined by
[…]
@@
for $a\in L(\mathcal{G})$. Lemma 3.4 remains valid, as does the Fourier decomposition of Proposition 3.5. We give the precise statement in the Proposition below but omit the proof since it carries over from Proposition 3.5 verbatim.
@@
Let $(\mathcal{G},\mu,\omega)$ be a discrete measured groupoid and fix a symmetric basis $\mathcal{B}$ of $\mathcal{G}$. Denote by $E:L_{\omega}(\mathcal{G})\to L^{\infty}(\mathcal{G}^{(0)},\mu)$ the conditional expectation. Any element of $a\in L_{\omega}(\mathcal{G})$ admits a Fourier decomposition
@@
First assume that $a\in\mathcal{Z}(L_{\omega}(\mathcal{G}))$. Fix $a\in\mathcal{Z}(L_{\omega}(\mathcal{G}))$. That $j(a)$ is supported on $\operatorname{Iso}(\mathcal{G})$ follows exactly as in the untwisted case. Take $g\in\mathcal{G}$ and $h\in\operatorname{Iso}(\mathcal{G})$ with $s(g)=s(h)$ and let $B\subset\mathcal{G}$ and $C\subset\operatorname{Iso}(\mathcal{G})$ be bisections containing $g$ and $h$ respectively. If $h\in\mathcal{G}^{(0)}$, take $C=\mathcal{G}^{(0)}$ and if $h\notin\mathcal{G}^{(0)}$, assume that $C\cap\mathcal{G}^{(0)}=\emptyset$. Fix a basis $\mathcal{B}$ of $\operatorname{Iso}(\mathcal{G})$ that contains $C$. On the one hand, we have
$j\left(\lambda_{B}a\lambda_{B^{-1}}\right)=\sum_{D\in\mathcal{B}}\omega(B,D)\omega(BD,B^{-1})(a^{D}\circ\sigma_{B}^{-1})\ast 1_{BDB^{-1}}.$
@@
$j\left(\lambda_{t(B)}a\right)=\sum_{D\in\mathcal{B}}a^{BDB^{-1}}\ast 1_{BDB^{-1}}.$
@@
Finally, note that whenever $k$ lies in some bisection $E$, on has that $j(a)(k)=a^{E}\ast 1_{E}(k)=a^{E}(t(k))$, from which (14) follows.
Conversely, let $B\subseteq\mathcal{G}$ be any bisection. We first show that $j(\lambda_{B}*a*\lambda_{B^{-1}})=j(\lambda_{t(B)}*a)$. For any $h\in\operatorname{Iso}(\mathcal{G})$, letting $g\in B$ be such that $s(g)=t(x)$ if $g$ exists,
$j(\lambda_{B}*a*\lambda_{B^{-1}})(h)$ $=e_{t(B)}(h)\omega(g^{-1},gh)\omega(g^{-1},h)j(a)(ghg^{-1})$
@@
$=j(\lambda_{t(B)}*a)(h)\;.$
Since $j$ is injective, we have the identity $\lambda_{B}*a*\lambda_{B^{-1}}=\lambda_{t(B)}*a$. Furthermore, since $j(a)$ is supported on $\operatorname{Iso}(\mathcal{G})$, it follows that for all $E\subseteq\mathcal{G}^{(0)}$, we have the identity $\lambda_{E}*a=a*\lambda_{E}$. Therefore,
$\lambda_{B}*a$ $=\lambda_{B}*\lambda_{s(B)}*a=\lambda_{B}*a*\lambda_{s(B)}$
$=\lambda_{t(B)}*a*\lambda_{B}=a*\lambda_{t(B)}*\lambda_{B}$
$=a*\lambda_{B}\;.$
Since the elements of the form $\lambda_{B}$ are SOT-dense in $L_{\omega}(\mathcal{G})$, we get the converse. ∎

### v3

@@ 3 Basis of a groupoid, Fourier decomposition, and the $j$ map
Since $a$ is central, $\lambda_{B}a\lambda_{B^{-1}}=\lambda_{B}\lambda_{B^{-1}}a=\lambda_{t(B)}a$. Looking at the base coordinate $C$, this in particular means that
@@
hence we get that $a^{BCB^{-1}}(t(ghg^{-1}))=a^{C}(t(h))$. Now it suffices to note that whenever an element $k\in\mathcal{G}$ lies in some bisection $E$, one has that $j(a)(k)=a^{E}\ast 1_{E}(k)=a^{E}(t(k))$. ∎
@@ 4 Factoriality of the untwisted groupoid von Neumann algebra
Let $(\mathcal{G},\mu)$ be an ergodic measured groupoid and let $A\subset\operatorname{Iso}(\mathcal{G})$ be a Borel subset with $\mu$-finite conjugacy class. Then there exist finitely many mutually disjoint bisections $V_{1},...,V_{k}$ of measure one such that $\Omega_{A}=\bigsqcup_{i}V_{i}$ up to a null set. In particular, $\mu_{s}(\Omega_{A})=k$.
@@
Let $q\colon\mathcal{R}_{\mathcal{G}}\rightarrow\mathcal{G}$ be a Borel section for the projection map $g\mapsto(t(g),s(g))$. Take an arbitrary $\varphi\in[\mathcal{R}_{\mathcal{G}}]$ and define $W=\{q((\varphi(x),x))\mid x\in\mathcal{G}^{(0)}\}$. Then $W\in[\mathcal{G}]$ and we have that $t(g)=\varphi(s(g))$ for every $g\in W$. Observe that
@@
By ergodicty we get that for each $k$ either $\mu(E_{k})=0$ or $\mu(E_{k})=1$. Since $\mathcal{G}^{(0)}=\bigsqcup_{k}E_{k}$, we find a $k$ for which $\mu(E_{k})=1$.
By [Kec95, Exercise 18.15], there exists a bisection $V_{1}\subseteq\Omega_{A}\cap s^{-1}(E_{k})$ such that $s(V_{1})=E_{k}$. Note that by definition of $E_{k}$, we have $s((\Omega_{A}\cap s^{-1}(E_{k}))\setminus V_{1})=E_{k}$. If $k=1$, then we are done. Otherwise, we work by induction: assume for some $j\leq k$ that we have mutually disjoint bisections $V_{1},\ldots,V_{j}\subseteq\Omega_{A}\cap s^{-1}(E_{k})$ with $s(V_{i})=E_{k}$ for all $i$. If $j=k$, then we would have the desired bisections. Otherwise, note that the map $s:(\Omega_{A}\cap s^{-1}(E_{k}))\setminus\left(\bigsqcup_{i\leq j}V_{i}\right)\to E_{k}$ must be a surjection by definition of $E_{k}$. Applying [Kec95, Exercise 18.15] again gives us a bisection $V_{j+1}\subseteq(\Omega_{A}\cap s^{-1}(E_{k}))\setminus\left(\bigsqcup_{i\leq j}V_{i}\right)$ with $s(V_{j+1})=E_{k}$. ∎
@@
When the ergodicity assumption in Lemma 4.4 is removed, it is straightforward to come up with a counterexample. The transposition $(1,2)$ has conjugacy class of size (${n\choose 2}$ in the symmetric group $\mathcal{S}_{n}$. Then writing $\rho=\sum_{n=2}^{\infty}2^{-n}{n\choose 2}^{-1}<+\infty$ we get that $\mu(n)=\rho^{-1}2^{-n}{n\choose 2}^{-1}$ defines a probability measure $\mu$ on $\mathbb{N}_{\geq 2}$. Define $\mathcal{G}=\{(n,\sigma)\mid n\geq 2\text{ and }\sigma\in\mathcal{S}_{n}\}$ with unit space $(\mathbb{N}_{\geq 2},\mu)$ and source and target map given by $(n,\sigma)\mapsto n$. The conjugacy class $\Omega_{A}$ of the bisection $A=\bigcup_{n\geq 2}\{(n,(1,2))\}$ satisfies $\mu(\Omega_{A})<+\infty$ by construction, but it cannot be written as the union of finitely many bisections.
@@
Now assume condition 2 fails. Then we find an element $a\in\mathcal{Z}(L(\mathcal{G}))\setminus L^{\infty}(\mathcal{G}^{(0)},\mu)$. Note that $j(a)\neq 0$ and we can assume that $j(a)|_{\mathcal{G}^{(0)}}=0$. Write
@@ 5.1 Definitions and basic properties
We say two cocycles $\omega_{1}$ and $\omega_{2}$ are cohomologous if there is a Borel map $\rho:\mathcal{G}\to\mathbb{T}$ such that $\omega_{1}(x,y)=\rho(x)\rho(y)\overline{\rho(xy)}\omega_{2}(x,y)$ for almost every $(x,y)\in\mathcal{G}^{(2)}$. Every cocycle is cohomologous to a normalized cocycle which moreover satisfies $\omega(x,x^{-1})=1$ for all $x\in\mathcal{G}$. To see this, first note that for any cocycle $\omega$ the cocycle identity implies that $\omega(x,x^{-1}x)=\omega(x^{-1}x,x^{-1}x)$ and $\omega(xx^{-1},x)=\omega(xx^{-1},xx^{-1})$ for all $x\in\mathcal{G}$. The cocycle
$\omega^{\prime}(x,y)=\overline{\omega(x,x^{-1}x)}\ \overline{\omega(y,y^{-1}y)}\omega(xy,(xy)^{-1}xy)\omega(x,y)$
satisfies $\omega^{\prime}(x^{-1}x,x^{-1}x)=1=\omega^{\prime}(xx^{-1},xx^{-1})$ for all $x\in\mathcal{G}$ and hence is normalized. Next, suppose $\omega$ is a normalized cocycle and consider
@@
Now we proceed analogously as in the untwisted case by defining a ‘projective left regular representation’ in the following way: for any Borel bisection $A\subset\mathcal{G}$ we define the operator $\lambda_{A}^{\omega}\in B(L^{2}(\mathcal{G},\mu_{s}))$ by
$\lambda_{A}^{\omega}1_{E}(g)=\omega(g_{1},g_{2})1_{A\cdot E}(g)\qquad\text{where }g_{1}\in A,g_{2}\in E\text{ such that }g_{1}g_{2}=g$
for any subset $E\subset\mathcal{G}$ with $\mu_{s}(E)<\infty$. The (left) twisted groupoid von Neumann algebra is defined as $L_{\omega}(\mathcal{G}):=\{\lambda_{A}^{\omega}\ |\ A\subset\mathcal{G}\text{ Borel bisection}\}^{\prime\prime}$. Note that, as in the untwisted case, $L_{\omega}(\mathcal{G})$ includes a copy of $L^{\infty}(\mathcal{G}^{(0)},\mu)$ as a regular subalgebra. For any Borel bisection $A$ one has $(\lambda_{A}^{\omega})^{*}=\lambda_{A^{-1}}^{\omega}$ and one can check that $\lambda_{A}^{\omega}\lambda_{B}^{\omega}=\omega(A,B)\lambda_{A\cdot B}^{\omega}$ for any two Borel bisections $A,B\subset\mathcal{G}$.
@@
$\rho_{A}^{\omega}1_{E}(g)=\overline{\omega(g_{2},g_{1})}1_{E\cdot A^{-1}}(g)\qquad\text{where }g_{1}\in A^{-1},g_{2}\in E\text{ such that }g_{2}g_{1}=g.$ (12)
One denotes the right twisted groupoid von Neumann algebra by $R_{\omega}(\mathcal{G}):=\{\rho_{A}^{\omega}\ |\ A\subset\mathcal{G}\text{ is a Borel bisection}\}^{\prime\prime}$. We again have the relations $(\rho_{A}^{\omega})^{*}=\rho_{A^{-1}}^{\omega}$ and $\rho_{A}^{\omega}\rho_{B}^{\omega}=\omega(A,B)\rho_{A\cdot B}^{\omega}$. We now show that the representations $\lambda^{\omega}$ and $\rho^{\overline{\omega}}$ are commutants of each other. For $f,g\in L^{2}(\mathcal{G},\mu_{s})$ we define the twisted convolution product $\ast_{\omega}$ by
@@
We say a function $f\in L^{2}(\mathcal{G})$ is a left convolver for $(\mathcal{G},\omega)$ if $f\ast_{\omega}g\in L^{2}(\mathcal{G},\mu_{s})$ for every $g\in L^{2}(\mathcal{G},\mu_{s})$. In that case it follows that $L_{f}^{\omega}(g)=f\ast_{\omega}g$ defines a bounded operator on $L^{2}(\mathcal{G},\mu_{s})$. We denote by $LC_{\omega}(\mathcal{G})$ the set consisting of all $L_{f}^{\omega}$ for left convolvers $f$. Similarly, one defines the right convolvers $RC_{\omega}(\mathcal{G})$.
@@
In particular as $f_{\mathcal{F}}\to f$ and $f_{cF}^{*}\to f^{*}$ in $L^{2}$-norm, the net $(\lambda_{f_{\mathcal{F}}}^{\omega})_{\mathcal{F}\in\mathscr{F}}$ is Cauchy. Since these elements belong in a fixed closed ball, $\lambda_{f_{\mathcal{F}}}^{\omega}$ converges in the $\sigma$-strong* operator topology to some $T\in L_{\omega}(\mathcal{G})$. We claim that $T=L_{f}^{\omega}$. For this, it suffices to show that these operators agree on bisections. For a bisection $B$, since the $\sigma$-strong* operator topology is stronger than the strong operator topology, we have the limit $T(1_{B})=\lim_{\mathcal{F}}\lambda_{f_{\mathcal{F}}}^{\omega}(1_{B})$, and
@@ 5.2 Fourier decomposition and the $j$ map again
still defines an injective linear map sending $\lambda_{A}^{\omega}$ to $1_{A}$ whenever $A\subset\mathcal{G}$ is a Borel bisection. As before, one can check that $j(\lambda_{A}^{\omega}x\lambda_{B}^{\omega})=\lambda_{A}^{\omega}\rho_{B^{-1}}^{\overline{\omega}}j(x)$. We again consider the faithful normal state $\phi_{\mu}$ on $L_{\omega}(\mathcal{G})$ uniquely determined by
$\phi_{\mu}(\lambda_{A})=\int E(1_{A})\mathrm{d}\mu=\mu(A\cap\mathcal{G}^{(0)})\text{ for every bisection }A\subset\mathcal{G}\,,$
@@
for $a\in L_{\omega}(\mathcal{G})$. Lemma 3.4 remains valid, as does the Fourier decomposition of Proposition 3.5. We give the precise statement in the proposition below but omit the proof since it carries over from Proposition 3.5 verbatim.
@@
Let $(\mathcal{G},\mu,\omega)$ be a twisted discrete measured groupoid and fix a symmetric basis $\mathcal{B}$ of $\mathcal{G}$. Denote by $E:L_{\omega}(\mathcal{G})\to L^{\infty}(\mathcal{G}^{(0)},\mu)$ the conditional expectation. Any element of $a\in L_{\omega}(\mathcal{G})$ admits a Fourier decomposition
@@
First assume that $a\in\mathcal{Z}(L_{\omega}(\mathcal{G}))$. That $j(a)$ is supported on $\operatorname{Iso}(\mathcal{G})$ follows exactly as in the untwisted case. Take $g\in\mathcal{G}$ and $h\in\operatorname{Iso}(\mathcal{G})$ with $s(g)=s(h)$ and let $B\subset\mathcal{G}$ and $C\subset\operatorname{Iso}(\mathcal{G})$ be bisections containing $g$ and $h$ respectively. If $h\in\mathcal{G}^{(0)}$, take $C=\mathcal{G}^{(0)}$ and if $h\notin\mathcal{G}^{(0)}$, assume that $C\cap\mathcal{G}^{(0)}=\emptyset$. Fix a basis $\mathcal{B}$ of $\operatorname{Iso}(\mathcal{G})$ that contains $C$. On the one hand, we have
$j\left(\lambda_{B}^{\omega}a\lambda_{B^{-1}}^{\omega}\right)=\rho_{B}^{\overline{\omega}}\lambda_{B}^{\omega}j(a)=\sum_{D\in\mathcal{B}}\omega(B,D)\omega(BD,B^{-1})(a^{D}\circ\sigma_{B}^{-1})\ast 1_{BDB^{-1}}.$
@@
$j\left(\lambda_{t(B)}^{\omega}a\right)=\sum_{D\in\mathcal{B}}a^{BDB^{-1}}\ast 1_{BDB^{-1}}.$
@@
Finally, note that whenever $k$ lies in some bisection $E$, one has that $j(a)(k)=a^{E}\ast 1_{E}(k)=a^{E}(t(k))$, from which (14) follows.
Conversely, let $B\subseteq\mathcal{G}$ be any bisection. We first show that $j(\lambda_{B}^{\omega}a\lambda_{B^{-1}}^{\omega})=j(\lambda_{t(B)}^{\omega}a)$. For any $h\in\operatorname{Iso}(\mathcal{G})$, letting $g\in B$ be such that $s(g)=t(h)$ if $g$ exists,
$j(\lambda_{B}^{\omega}a\lambda_{B^{-1}}^{\omega})(h)$ $=1_{t(B)}(h)\omega(g^{-1},gh)\omega(g^{-1},h)j(a)(ghg^{-1})$
@@
$=j(\lambda_{t(B)}^{\omega}a)(h)\;.$
Since $j$ is injective, we have the identity $\lambda_{B}^{\omega}a\lambda_{B^{-1}}^{\omega}=\lambda_{t(B)}^{\omega}a$. Furthermore, since $j(a)$ is supported on $\operatorname{Iso}(\mathcal{G})$, it follows that for all $E\subseteq\mathcal{G}^{(0)}$, we have the identity $\lambda_{E}^{\omega}a=a\lambda_{E}^{\omega}$. Therefore,
$\lambda_{B}^{\omega}a$ $=\lambda_{B}^{\omega}\lambda_{s(B)}^{\omega}a=\lambda_{B}^{\omega}a\lambda_{s(B)}^{\omega}$
$=\lambda_{t(B)}^{\omega}a\lambda_{B}^{\omega}=a\lambda_{t(B)}^{\omega}\lambda_{B}^{\omega}$
$=a\lambda_{B}^{\omega}\;.$
Since the elements of the form $\lambda_{B}^{\omega}$ are SOT-dense in $L_{\omega}(\mathcal{G})$, we get the converse. ∎